Reports until 13:41, Monday 31 March 2025
GQuEST Output Filter Cavities
Torrey Cullen - posted 13:41, Monday 31 March 2025 (12204)
OFC3 Initial 775 Lock

Continuation of 12198.

I removed the lens in transmission of the cavity. This allowed me to see it was cleary clipping something circular. I walked the input beam and intracavity mirrors until the clipping was no longer observed. With this clipping gone was able to tune the alignment by eye on the camera a little more. At this point it is much better to use PDs to fine tune the alignment. I installed a 50:50 beam splitter in transmission as well as a PD. Additionally, I installed 2 mirrors in reflection of the cavity that act as steering mirrors onto a REFL PD. Next I aligned onto the REFL PD to allow for locking. While the cavity is locked, you can see the TRANS beam on the card to allow for easier alignment onto the TRANS PD.

775 path is fully installed and aligned. A cavity scan shows the cavity is well aligned. While locked, it seems there is some periodic signal on the error signal (this is laser locked with the teraxion) that is causing some drops in the amount of transmitted light power. By changing the coupling on REFL to DC you can check the visibility on the REFL PD. It seems pretty good, but may be improved by a lens as the diode is very small on these fast PDs used for PDH.

Images attached to this report